OpenStudy (anonymous):

x+y=10 what is the x and y intercept

OpenStudy (alexwee123):

to find the x intercept you would make y=0 then solve for x to find the y intercept you would make x=0 then solve for y

OpenStudy (anonymous):

thank youu :)

OpenStudy (anonymous):

how would i do 9x=54-6y?

OpenStudy (anonymous):

set it equal to 0 and solve for the variable..

OpenStudy (anonymous):

9x=54-6y 9x=54-6(0) 9X=54-6 9X/9=48/9

OpenStudy (anonymous):

5.3?

OpenStudy (anonymous):

so set x to 0 and solve for y for the y intercept.. if its easier for you to see, you can rewrite the equation.

OpenStudy (anonymous):

9x + 6y = 54

OpenStudy (anonymous):

now.. substitute 0 in for x.. so 9(0) +6y = 54... so 6y = 54 .. so y = 54 / 6 ... thus y = 9.. so the y intercept is (0,9)

OpenStudy (anonymous):

now take your same equation and plug in 0 for y making the equation 9x + 6(0) = 54. can you get it from here?

OpenStudy (anonymous):

Ohh so you have to put it in standard form first?

OpenStudy (anonymous):

you can do it without I believe.. but yes, its easier if you move all variable terms to one side and the constant terms to the other side, at least when doing the intercepts.

OpenStudy (anonymous):

okay, thanks so much!!

OpenStudy (anonymous):

no problem!! best of luck!! can you medal my answer if I helped you please? :)

OpenStudy (anonymous):

yep ! :)